# Layers at Similar Depths Generate Similar Activations Across LLM Architectures

### v1

@@ 5 Conclusion
We compare the activations of a wide range of modern LLMs and find that they generate different activation geometries at different depths (Claim 1), but that layers at corresponding depths from different models tend to generate similar activation geometries (Claim 2). We find that early layers from different models are just as similar to one another as middle and late layers, and validate our findings by developing a statistical test for diagonal structure.
@@ Appendix A Diagonal structure is sensitive to the input distribution
The experiments above all focused on the similarity of nearest neighbor relationships on a dataset of inputs from OpenWebText. In Appendix E, we generate affinity matrices for a range of input datasets, including IMDB movie reviews (Maas et al., 2011), text from books in English and German (Tiedemann, 2012), and various LLM evaluations (Zhou et al., 2023; Hendrycks et al., 2021). The diagonal structure observed in Section 4 is generally present for other distributions of web text (e.g., IMDB movie reviews), but different patterns emerge when looking at other input distributions, which we investigate here.
@@ A.1 Nearest neighbor relationships are (weakly) preserved between languages
Many of the models we use in our experiments have been trained on multilingual data, and can write and understand multiple languages. In the experiments above, we feed the same dataset $\mathcal{D}$ to two models $M_{1}$ and $M_{2}$ and collect the resulting activations. Here, using a parallel corpus of English and German book translations (Tiedemann, 2012), we instead feed a corpus of English texts to $M_{1}$ and a parallel corpus of translated German texts to $M_{2}$. The resulting nearest neighbor sets exhibit lower similarity than when using only a single language, but we still observe weak diagonal structure, with middle layers being more similar to one another than to early or late layers (Fig. 7).
@@ A.2 Instruction tuning changes activation structure in late layers
Instruction tuning is a process in which a base model (which has generally been pretrained on natural language) is fine-tuned to be better at following instructions (Zhang et al., 2023). We investigate how instruction tuning affects the activations generated by language models.
We compare the base version of Gemma-2-9B to its instruction-tuned counterpart. When given inputs from OpenWebText, we observe strong diagonal structure across all depths (Fig. 8(a)). However, when given prompts from a dataset of inputs designed to test a model’s ability to follow instructions (Zhou et al., 2023) there is much lower similarity at later layers (Fig. 8(b)).
Figure 8: Base models and their instruction-tuned counterparts generate similar activations at corresponding depths when given web text, but not when given instruction . Instructing tuning changes the activations generated by models in later layers, but only when given inputs that are targed by instruction tuning. 8(a) The affinity matrix of the base version of Gemma-2-9B and its instruction tuned counterparts, given a dataset of inputs from OpenWebText. 8(b) The same as 8(a), but using a dataset of inputs from IFEval.
We hypothesize that later layers could be dedicated to functions that are strongly targeted by instruction tuning, while early layers could be dedicated to functions like input parsing and basic grammar that are needed with or without instruction tuning. Alternatively, the instructing tuning process used could have been designed to only train later layers.
@@ A.3 Random alphanumeric strings
Fig. 9 reveals that there is similarity between the nearest neighbor relationships of random strings, but without diagonal structure. Just like with web text in Section 4, layers from different models often agree on the nearest neighbor relationships among the random strings. However, unlike with web text, the most similar layers do not necessarily come from corresponding depths. Instead, all early and late layers show similarity with one another, while middle layers are dissimilar between models.
@@
Appendix B Null model for mutual k-nearest neighbors
@@
$M^{(\mathcal{D})}_{k}(f,g)$ $=\frac{1}{k\|\mathcal{D}\|}\sum_{x\in\mathcal{D}}\left\|\mathcal{U}^{(\mathcal{D})}_{k}(x,f)\cap\mathcal{U}^{(\mathcal{D})}_{k}(x,g)\right\|$ (5)
$\sim\mathrm{Normal}\left(\frac{k}{|D|-1},\sqrt{\frac{(k-\|\mathcal{D}\|+1)^{2}}{\|\mathcal{D}\|(\|\mathcal{D}\|-2)(\|\mathcal{D}\|-1)^{2}}}\right)$ (6)
which gives the distribution of $M^{(\mathcal{D})}_{k}(f,g)$ under the null assumption made above.
Figure 10: Null distribution of $M^{(\mathcal{D})}_{k}(f,g)$ with $k=10$ and $\|\mathcal{D}\|=2048$. The log pdf is astronomically small for values larger than approximately $0.02$.
@@
The null distribution derived above is extremely concentrated, with the null probability of $M^{(\mathcal{D})}_{k}(f,g)\geq 0.4$ being on the order of $10^{-143,449}$. Thus, the mutual $k$-NN similarities scores appearing in these affinity matrices are far higher than would be expected under this (very strong) null model.
@@
Appendix C Model list
@@
Appendix D Other values of k
In the experiments above, we measure similarity with mutual $k$-NN with $k=10$. Here, we generate versions of Fig. 4(a) using various other values of $k$ (Fig. 11). The results show similar diagonal structure, even with $k=1$ (which corresponds to the probability that the layers being compared agree on the single nearest neighbor).
Appendix E Extended affinity matrix arrays
We computed affinity matrices comparing all pairs of all layers in 24 open-weight LLMs. For the sake of brevity, we only included a subset in Fig. 4(c). This appendix contains affinity matrices for all models on all datasets that we tested.
E.1 Rectangular affinity matrices
@@
E.2 Square affinity matrices
These affinity matrices have been stretched and squished to fit into squares to make them more comparable, as in Fig. 4(c).

### v3

@@ 5 Conclusion
We compare the activations of a wide range of modern LLMs and find that they generate different activation geometries at different depths (Claim 1), but that layers at corresponding depths from different models tend to generate similar activation geometries (Claim 2). We find that early layers from different models are just as similar to one another as middle and late layers, and validate our findings by developing a statistical test of diagonal structure.
Acknowledgments
We would like to thank Claire Donnat and Nikos Ignatiadis for their advice on statistical tests of diagonal structure.
@@ Appendix A Diagonal structure is sensitive to the input distribution
The experiments above all focused on the similarity of nearest neighbor relationships on a dataset of inputs from OpenWebText. In Appendix G, we generate affinity matrices for a range of input datasets, including IMDB movie reviews (Maas et al., 2011), text from books in English and German (Tiedemann, 2012), and various LLM evaluations (Zhou et al., 2023; Hendrycks et al., 2021). The diagonal structure observed in Section 4 is generally present for other distributions of web text (e.g., IMDB movie reviews), but different patterns emerge when looking at other input distributions, which we investigate here.
@@ A.1 Nearest neighbor relationships are (weakly) preserved between languages
Many of the models we use in our experiments have been trained on multilingual data, and can write and understand multiple languages. In the experiments above, we feed the same dataset $\mathcal{D}$ to two models $M_{1}$ and $M_{2}$ and collect the resulting activations. Here, using a parallel corpus of English and German book translations (Tiedemann, 2012), we instead feed a corpus of English texts to $M_{1}$ and a parallel corpus of translated German texts to $M_{2}$. The resulting nearest neighbor sets exhibit lower similarity than when using only a single language, but we still observe weak diagonal structure, with middle layers being more similar to one another than to early or late layers (Figure 7).
@@ A.2 Instruction tuning changes activation structure in late layers
Instruction tuning is a process in which a base model (which has generally been pretrained on natural language) is fine-tuned to be better at following instructions (Zhang et al., 2024). We investigate how instruction tuning affects the activations generated by language models.
We compare the base version of Gemma-2-9B to its instruction-tuned counterpart. When given inputs from OpenWebText, we observe strong diagonal structure across all depths (Figure 8(a)). However, when given prompts from a dataset of inputs designed to test a model’s ability to follow instructions (Zhou et al., 2023) there is much lower similarity at later layers (Figure 8(b)). This finding is in line with previous work that observed later layers being more strongly affected by fine-tuning (Wu et al., 2020).
Figure 8: Instruction tuning changes the activations generated by models in later layers, but only when given inputs that are targeted by instruction tuning. 8(a) The affinity matrix of the base version of Gemma-2-9B and its instruction tuned counterparts, given a dataset of inputs from OpenWebText. 8(b) The same as 8(a), but using a dataset of inputs from IFEval.
We hypothesize that later layers could be dedicated to functions that are strongly targeted by instruction tuning, while early layers could be dedicated to functions like input parsing and basic grammar that are needed with or without instruction tuning. Alternatively, the instruction tuning process used could have been designed to only train later layers.
@@ A.3 Random alphanumeric strings
Figure 9 reveals that there is similarity between the nearest neighbor relationships of random strings, but without diagonal structure. Just like with web text in Section 4, layers from different models often agree on the nearest neighbor relationships among the random strings. However, unlike with web text, the most similar layers do not necessarily come from corresponding depths. Instead, all early and late layers show similarity with one another, while middle layers are dissimilar between models.
@@
Appendix B Block bootstrap test of diagonal structure
Given an affinity matrix, we want to test whether the on-diagonal cells show higher similarity scores than the off-diagonal cells. Naively, we could perform a one-sided $t$-test comparing the scores in on- and off-diagonal cells, but this would assume that each cell is independent, which need not be true because of residual connections in transformer models (Vaswani et al., 2017).
(a) An affinity matrix AA
(b) Bootstrap samples where AA has been scrambled with a block size of $10\times 10$.
Figure 10:
We use the block bootstrap to generate samples from a null distribution that preserves local dependence structure, but scrambles large-scale structure (Lahiri, 1999; Davison & Hinkley, 1999). The block bootstrap works by breaking an affinity matrix into blocks of a given size and then resampling those blocks (with replacement) to create a new matrix that inherits the local dependence of the original, but does not inherit its global structure (Figure 10). We use overlapping blocks with a cyclic boundary condition so that every cell has an equal probability of appearing in the resampled matrix. By comparing the difference of on- and off-diagonal means in the observed affinity matrices with their corresponding resampled matrices, we test whether the observed diagonal structure is stronger than would be expected as a consequence of the local dependence structure and chance.
For each affinity matrix, we generate $10^{5}$ scrambled bootstrap samples and compute the difference of the on- and off-diagonal means. The fraction of bootstrap samples that yields a difference at least as large as that observed from the original affinity matrix gives us a $p$-value (Figure 11).
Figure 11: Distribution of differences between on- and off-diagonal means for bootstrap samples for a particular affinity matrix. The red line shows the difference between on- and off-diagonal means observed with the unscrambled affinity matrix. In Figure 11(b), the blue line shows 0.95.
We used this process to compute a $p$-value for every affinity matrix (generated on OpenWebText data), using block sizes of $5\times 5$ and $10\times 10$. For most model pairs, there was not a single bootstrap sample with an on- and off-diagonal difference as large as the observed affinity matrix (Figure 12). No model pair yielded a $p$-value greater than 0.05 with either block size. The maximum $p$-value with block size $5\times 5$ is 0.00484, and the maximum $p$-value with block size $10\times 10$ is 0.0154.
Figure 12: Every test yields a $p$-value below 0.05. 12(a) Manhattan plot generated with a block size of $5\times 5$. Tests where no bootstrap sample had a difference as great as the observed difference are clipped to one over the number of bootstrap samples for visualization purposes. 12(b) Manhattan plot generated with a block size of $10\times 10$.
We can combine these tests into a single intersection-union test of whether the on-diagonal similarity is higher than the off-diagonal similarity for every pair of models simultaneously. Because we can reject the null-hypothesis for each of the individual tests (as the maximum $p$-value is below 0.05), we can reject it for the intersection-union test, giving us the strong result that every pair of models shows significant diagonal structure (Berger, 1982; Kim & Allison, 2006).
Appendix C Null model for mutual k-nearest neighbors
@@
$\mathcal{M}^{{(\mathcal{D})}}_{k}(f,g)$ $=\frac{1}{k\|\mathcal{D}\|}\sum_{x\in\mathcal{D}}\left\|\mathcal{U}^{{(\mathcal{D})}}_{k}(x,f)\cap\mathcal{U}^{{(\mathcal{D})}}_{k}(x,g)\right\|$ (5)
$\sim\mathrm{Normal}\left(\frac{k}{\|D\|-1},\sqrt{\frac{(k-\|\mathcal{D}\|+1)^{2}}{\|\mathcal{D}\|(\|\mathcal{D}\|-2)(\|\mathcal{D}\|-1)^{2}}}\right)$ (6)
which gives the distribution of $\mathcal{M}^{{(\mathcal{D})}}_{k}(f,g)$ under the null assumption made above.
Figure 13: Null distribution of $\mathcal{M}^{{(\mathcal{D})}}_{k}(f,g)$ with $k=10$ and $\|\mathcal{D}\|=2048$. The log pdf is astronomically small for values larger than approximately $0.02$.
@@
The null distribution derived above is extremely concentrated, with the null probability of $\mathcal{M}^{{(\mathcal{D})}}_{k}(f,g)\geq 0.4$ being on the order of $10^{-143{,}449}$ (when $k=10$ and $\|\mathcal{D}\|=2048$). Thus, the mutual $k$-NN similarity scores appearing in the affinity matrices above are far higher than would be expected under this (very strong) null model.
@@
Appendix D Other values of k
In the experiments above, we measure similarity with mutual $k$-NN with $k=10$. Here, we generate versions of Figure 4(a) using various other values of $k$ (Figure 14). The results show similar diagonal structure, even with $k=1$ (which corresponds to the probability that the layers being compared agree on the single nearest neighbor).
Appendix E Decay curves
The following plots visualize how the relative depth of layers relates to their similarity. In particular, they show how similarity decays as distance increases.
Figure 15: Nearby layers within a model tend to be more similar than distant ones. For a given model, each plot shows the similarity between layers as a function of their distance. Each line corresponds to a layer with the color indicating its depth. The $x$-axis shows the relative depth of another layer in the same model, and the $y$-axis shows their similarity. Each curve can be thought of as showing a slice of the affinity matrices above. All plots use mutual $k$-nearest neighbors with OpenWebText as the dataset.
Figure 16: Corresponding layers from different models tend to be more similar than non-corresponding ones. For a given pair of models, each plot shows the similarity between layers in one model to layers in another model as those layers get further away from the path of maximum similarity. Each line corresponds to a layer in one model with the color indicating its depth. The $x$-axis shows the depth of the layer in the other model, shifted so that the highest-similarity layer is at 0. The $y$-axis shows the similarity of the two layers. Each curve can be thought of as showing a slice of the affinity matrices above. All plots use mutual $k$-nearest neighbors with OpenWebText as the dataset.
Appendix F Model list
@@
Appendix G Extended affinity matrix arrays
We computed affinity matrices comparing all pairs of all layers in 24 open-weight LLMs. For the sake of brevity, we only included a subset in Figure 4(c). This appendix contains affinity matrices for all models on all datasets that we tested.
G.1 Rectangular affinity matrices
@@
G.2 Square affinity matrices
These affinity matrices have been stretched and squished to fit into squares to make them more comparable, as in Figure 4(c).